one of two complementary angles is 16 degrees less than its complement. find the mearure of both angles.
Answer by arKed(27)

N1 +N2 = 90 eq.no1
N1 = N2 – 16 eq. no2
N2 = N1 = 16 eq. no3
Sol’n:
1.) Substitute equation no. 3 to eq. no. 1
N1 + N2 = 90
N2 – 16 + N2 = 90
2N2 = 90 +16
2N2 = 106
2N2 / 2 = 106 / 2
N2 = 53 ® ANS.

2.) Substitute N2 = 53 ® to eq. no. 2
N1 = N2 -16
N1 = 53 – 16
N1 = 37® ANS.